# LLM-Assisted Request Mutation for Whitebox REST API Testing

### v2

@@ 4.4. RQ3: Runtime Overhead
SUT Num of Mutation Times Average Execution Time per Test (ms)
@@
To evaluate the runtime overhead introduced by LLM-assisted mutation, we use Number of Mutation Times as a metric. Since the duration of each test is fixed at 1 hour, this metric can measure the number of mutations performed within the same time under different experimental settings. By comparing the number of mutations, we can assess the runtime overhead: the more mutations performed within the same time, the lower the time cost per mutation, indicating a smaller runtime overhead. We also calculate Average Execution Time per Test to directly measure the time cost of every run of test case mutation and evaluation. Table 4 shows the results in detail for the runtime overhead between three different configurations: the baseline, our proposed method, and our method without variable expansion (no VE).
It shows that our method does cause more runtime overhead than baseline, but not more. Our method got an 18% deduction of mutation times from the baseline. In terms of average execution time per test, our method generally results in longer execution times compared to the baseline, which aligns with the intuition that the query of a language model (LLM) is expected to incur additional runtime overhead. However, the increase in runtime overhead is relatively modest, amounting to only 31% of the baseline.
In summary, although our method increases runtime overhead, this increase is entirely acceptable given the substantial gains in accuracy. Despite the reduction in the number of mutations, the enhanced accuracy of our method results in an overall increase in line coverage.
@@ 5. Threats to validity
Internal. One threat to internal validity in our work is from the randomness of MIO and LLMs. To address randomness, we conducted repeated experiments three times. Another threat comes from the dependency of the underlying search algorithm. We preserve the target selection strategy in the MIO algorithm, which may not be the most adaptive for our LLM-assisted mutation. In our experiments, we found that four benchmark programs were excluded due to the target selection process resulting in no targets within the scope of analysis. In future work, we plan to co-design with other components inside the MIO algorithm to achieve the best overall performance.
External. The generalizability of our tool’s testing ability might be limited by several factors. One comes from our evaluations only focus on open-source services written in Java language and utilize specific language models GPT-4o. The risk of generalizability across different programming languages can be resolved by the same comprehension ability of ChatGPT across different programming languages (e.g., Java, JavaScript, Python, and C) (Qiu et al., 2024). Besides, def-use analysis is a common static analysis that can be applied to other languages. These indicate that our method is compatible with different languages. Another risk comes from the metric we select. Although there is a very strong correlation between the coverage achieved and the number of bugs found by a fuzzer, fuzzers best at achieving coverage, may not be best at finding bugs (Böhme et al., 2022). Nevertheless, given that coverage is the most widely used metric, we continue to employ it. As for benchmark generalizability, we have included various REST API services of different sizes and functionality to ensure generalizability. However, the effectiveness of our method on industrial-level services has not been empirically verified, which is a limitation of our current evaluation. Additionally, the requirement for source code access may limit the applicability of our tool in scenarios where source code is not available, thus affecting its generalizability to closed-source or proprietary software.
@@ 6. Related Work
Autamoated API Testing. Blackbox API testing infers inter-
@@
parameter dependencies by analyzing API schemas. RESTler (Atlidakis et al., 2019) generates the correct sequence of API requests by producer-consumer dependencies. Morest (Liu et al., 2022) generate RESTful Service Property Graph (RPG) using extracted dependencies between APIs. Besides, there are other works resolving oracle problem (Segura et al., 2018; Luo et al., 2019), handling database (Zhang and Arcuri, 2021) and so on. Whitebox API testing focuses on the design of heuristics (Arcuri, 2017), and to solve the common issue in SBST, the flag problem (Baresel and Sthamer, 2003), testability transformations (Arcuri and Galeotti, 2021) are proposed to transform the SUT’s source code in a way that enhances the fitness function. Our work is orthogonal to the enhancement of fitness function, as our focus is on improving the accuracy of mutation operations. Therefore, these two aspects can complement each other in a synergistic manner.
@@
LLMs for Unit Test Generation. Current efforts to enhance the effectiveness of unit test generation with LLMs encompass several strategies. These include pre-training or fine-tuning of LLMs (Alagarsamy et al., 2023; Shin et al., 2024; Khanfir et al., 2023), the design of effective prompts using focal context (Chen et al., 2024), including focal method and focal class or continuous improvement by a generator and refiner (Yuan et al., 2024), and test generation with additional documentation, which incorporates API documentation (Vikram et al., 2024) and user-written bug reports (Plein et al., 2023). Furthermore, the integration of LLMs with traditional Search-Based Software Testing (SBST) is being explored (Lemieux et al., 2023). In instances where SBST stalls, the LLM can be queried for sample test cases.
@@
LLMs for Repository-Level Task. Repository-Level tasks, including package migration (Bairi et al., 2023), issue resolution (Ma et al., 2024; Xia et al., 2024; Zhang et al., 2024), and code completion (Phan et al., 2024; Liu et al., 2024b), necessitate a thorough analysis of the entire code repository. The objective of this analysis is to excavate and understand the intricate web of dependencies that exist within the repository. This broader perspective allows for a more holistic understanding of the codebase, enabling more effective and efficient solutions to the tasks at hand. CodePlan represents these dependencies as a dependency graph, encompassing syntactic relations, import relations, inheritance relations, method override relations, method invocation relations, object instantiation relations, and field use relations. RepoUnderstander delves into these dependencies by constructing a repository knowledge graph. Agentless, on the other hand, employs the repository structure format to construct a succinct representation of the repository’s file and directory structure. RepoHyper constructs the Repo-level Semantic Graph (RSG), a novel semantic graph structure that encapsulates the vast global context of code repositories at the method level. This graph includes import relations, invoke relations, ownership relations, encapsulate relations, and class hierarchy relations. GraphCoder enhances the accuracy of capturing the context of the completion target through a code context graph (CCG), which comprises control-flow, data-, and control dependence at the statement level.
@@ 7. Conclusion
We present MioHint, an LLM-assisted request mutator designed to address the fitness plateau issue associated with inefficient random mutation. MioHint generates precise mutation hints using LLMs with its code comprehension ability. To identify the relationship between requests and target conditions, MioHint performs cross-file, cross-procedure data flow analysis to gather global context for the target, which enables more accurate targeted mutation guidance. Evaluation results on 16 real-world REST API services across variant functionalities demonstrate that MioHint outperforms the state-of-the-art white-box API testing tool EvoMaster by 67$\times$ in mutation accuracy and achieves 4.95% higher line coverage. Furthermore, our method is able tp covers over 57% of hard-to-cover targets, whereas the baseline achieves coverage of less than 10%.

### v3

@@ 4.4. RQ3: Runtime Efficiency
SUT # Mutations Avg. Time per Test (ms)
@@ 5. Threats to Validity
Internal. A threat to the internal validity of our work arises from the randomness inherent in MIO and LLMs. To mitigate this threat, we repeated each experiment 10 times. Another threat relates to our dependency on the underlying search algorithm. Specifically, we retained the original target selection strategy of the MIO algorithm, which may not be optimal for our LLM-assisted mutation approach. In our experiments, this led to the exclusion of four benchmark programs because the process failed to identify any targets within the designated analysis scope. For future work, we plan to co-design our mutation strategy with other components of the MIO algorithm to improve overall performance.
External. The generalizability of our tool’s testing ability might be limited by several factors. One comes from our evaluations, which only focus on open-source services written in Java and utilize specific language models, GPT-4o. The risk of generalizability across different programming languages can be resolved by the same comprehension ability of ChatGPT across different programming languages (e.g., Java, JavaScript, Python, and C) (Qiu et al., 2024). Besides, def-use analysis is a common static analysis that can be applied to other languages. These indicate that our method is compatible with different languages. Another risk comes from the metric we select. Although there is a very strong correlation between the coverage achieved and the number of bugs found by a fuzzer, fuzzers best at achieving coverage, may not be best at finding bugs (Böhme et al., 2022). Nevertheless, given that coverage is the most widely used metric, we continue to employ it. Additionally, the requirement for source code access may limit the applicability of our tool in scenarios where source code is not available, thus affecting its generalizability to closed-source or proprietary software.
@@ 6. Related Work
Automated API Testing. Blackbox API testing analyzes API specifications to infer dependencies between parameters. RESTler (Atlidakis et al., 2019) generates the correct sequence of API requests by producer-consumer dependencies. Morest (Liu et al., 2022) generates a RESTful Service Property Graph (RPG) using extracted dependencies between APIs. Besides, there are other works resolving oracle problem (Segura et al., 2018; Luo et al., 2019), handling database (Zhang and Arcuri, 2021), and so on. Recent advances have explored using LLMs to further improve blackbox API testing, including: API specifications enhancement by identifying rules and generating example parameter values using LLM (Kim et al., 2024a); automated specifications inference using LLM for black-box testing even without realistic specification (Decrop et al., 2024); and multi-agent reinforcement learning that incorporates a semantic property dependency graph alongside LLMs to simplify the search space for dependencies and optimize API exploration (Kim et al., 2024b). Whitebox API testing focuses on the design of heuristics (Arcuri, 2017). To solve the common issue in SBST, the flag problem (Baresel and Sthamer, 2003), testability transformations (Arcuri and Galeotti, 2021) are proposed to transform the SUT’s source code in a way that enhances the fitness function. Our work is orthogonal to the enhancement of fitness function, as our focus is on improving the accuracy of mutation operations. Therefore, these two aspects can complement each other in a synergistic manner.
@@
LLMs for Unit Test Generation. Current efforts to enhance the effectiveness of unit test generation with LLMs encompass several strategies. These include pre-training or fine-tuning of LLMs (Alagarsamy et al., 2024; Shin et al., 2024; Khanfir et al., 2023), the design of effective prompts using focal context (Chen et al., 2024), including focal method and focal class or continuous improvement by a generator and refiner (Yuan et al., 2024), and test generation with additional documentation, which incorporates API documentation (Vikram et al., 2024) and user-written bug reports (Plein et al., 2023). Furthermore, the integration of LLMs with traditional Search-Based Software Testing (SBST) is being explored (Lemieux et al., 2023). In instances where SBST stalls, the LLM can be queried for sample test cases. These approaches are fundamentally limited by their narrow scope. Designed for unit testing, their analysis is inherently confined to isolated code snippets like individual functions. In contrast, our work targets system-level API testing, which demands a holistic understanding of interactions spanning the entire repository. Consequently, these unit-level techniques are insufficient as they overlook the broader codebase context required to understand the critical cross-module dependencies and end-to-end behaviors.
@@
LLMs for Repository-Level Task. Repository-Level tasks, including package migration (Bairi et al., 2023), issue resolution (Ma et al., 2024; Xia et al., 2024; Zhang et al., 2024), and code completion (Phan et al., 2024; Liu et al., 2024b), necessitate a thorough analysis of the entire code repository. The objective of this analysis is to excavate and understand the intricate web of dependencies that exist within the repository. This broader perspective allows for a more holistic understanding of the codebase, enabling more effective and efficient solutions to the tasks at hand. CodePlan represents these dependencies as a dependency graph, encompassing syntactic relations, import relations, inheritance relations, method override relations, method invocation relations, object instantiation relations, and field use relations. RepoUnderstander delves into these dependencies by constructing a repository knowledge graph. Agentless, on the other hand, employs the repository structure format to construct a succinct representation of the repository’s file and directory structure. RepoHyper constructs the Repo-level Semantic Graph (RSG), a novel semantic graph structure that encapsulates the vast global context of code repositories at the method level. This graph includes import relations, invoke relations, ownership relations, encapsulation relations, and class hierarchy relations.
While repository-level approaches recognize the importance of a global perspective, they differ in granularity and purpose. Typically, they build dependency or semantic graphs at the method or class level, resulting in coarse representations that are suitable for tasks like code completion or migration but inadequate for system-level testing. Such methods often include irrelevant code, obscuring the critical value transmission path from request to target. In contrast, our approach conducts statement-level data flow analysis, enabling the extraction of precise and minimal code slices that accurately capture the value propagation of an API call. This fine-grained context is essential for effective test generation and targeted mutation, offering a level of precision not achieved by existing techniques.
@@ 7. Conclusion
[…]